OpenStudy (anonymous):

The nucleus of the helium atom contains two protons that are separated about 3.0x10^-15m. Find the magnitude of the electrostatic force between the protons. (The protons remains together in the nucleus because the repulsive electrostatic force is balanced by an attractive force called the string nuclear force.)

OpenStudy (anonymous):

F=(kq^2)/(r^2) substitute the values. q=charge of proton

OpenStudy (anonymous):

What happened to the other \[q_{2}\]

OpenStudy (anonymous):

q1 and q2 are both protons, hence they have equal charge.

OpenStudy (anonymous):

then what what is the number for protons ?

OpenStudy (anonymous):

you mean the charge carried by a proton?

OpenStudy (anonymous):

I know k= 8.99x10^9

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

the charge of a proton is 1.60217646 × 10-19 coulombs

OpenStudy (anonymous):

\[q1q2\div4pi \epsilon0 r ^{2}\] So squaring the distance gives 9^-30 multiply by 4 pi epsilon zero. That gives the denominator. So the bottom of the quotient will be very small (1.01^-39) The top is the charge squared which is 1.602^-19 x 1.602^-19 which gives 2.566^-38. So the force is the top divided by the bottom which is 2.566^-38/1.01^-39= 25 N.

OpenStudy (anonymous):

Okay thank you . Very helpful